OpenStudy (anonymous):

A solution is made by dissolving 25.5 grams of glucose (C6H12O6) in 398 grams of water. What is the freezing-point depression of the solvent if the freezing point constant is -1.86 °C/m? Show all of the work needed to solve this problem.

OpenStudy (jfraser):

turn the mass of glucose into moles, and plug into the equation for freezing-point depression\[\Delta T_f = (\frac{mol solute}{kg solvent})* K_f\]

OpenStudy (anonymous):

can you show me how to do it ? its been haunting me all day

OpenStudy (jfraser):

turn mass into moles. use the molar mass. If you're doing FP depression, you know how to turn mass into moles

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can you walk me through it?

OpenStudy (jfraser):

moles = mass divided by molar mass: \[moles = mass * (\frac{moles}{mass})\]

OpenStudy (anonymous):

can you fill it in and i can try the math, and you check it?

OpenStudy (jfraser):

i've already done the math, you tell me what you get ;-)

OpenStudy (anonymous):

1 mol, don't know what the mass is

OpenStudy (jfraser):

you've got the formula of glucose, find the molar mass using the PT

OpenStudy (anonymous):

180.09

OpenStudy (jfraser):

that's the molar mass, now find the moles of 25.5g

OpenStudy (anonymous):

how? lol :/

OpenStudy (anonymous):

?

OpenStudy (jfraser):

plug the mass and molar mass into the formula i posted above

OpenStudy (anonymous):

i dont know how. can you do it? only plug it in tho

OpenStudy (jfraser):

the mass given is 25.5g. the molar mass is 180g.

OpenStudy (anonymous):

25.5*1/180

OpenStudy (jfraser):

yes, but that's just finding moles. those moles have to go into the equation i posted first \[\Delta T_f = (\frac{mol solute}{kg solvent})* K_f\]\[\Delta T_f = (\frac{0.142mol solute}{kg solvent})* K_f\] now find the rest and plug in to find DT

OpenStudy (anonymous):

whats the * k

OpenStudy (jfraser):

Kf is the freezing point depression constant, and it's given

OpenStudy (anonymous):

is the kg of solvent given?

OpenStudy (jfraser):

the grams of solvent are given, turn it into kg

OpenStudy (anonymous):

C6H12O6 is this the solvent

OpenStudy (anonymous):

.398 kg

OpenStudy (jfraser):

the glucose is the solUTE that you just solved for. what's the solVENT?

OpenStudy (anonymous):

water

OpenStudy (jfraser):

the 398g of water is the solvent now plug in the proper values and solve

OpenStudy (anonymous):

-.663

OpenStudy (anonymous):

correct?

OpenStudy (jfraser):

that's what i get

OpenStudy (anonymous):

can you check this?

OpenStudy (anonymous):

A solution is made by dissolving 4.0 moles of sodium chloride (NaCl) in 2.05 kilograms of water. If the molal boiling point constant for water (Kb) is 0.51 °C/m, what would be the boiling point of this solution? Show all of the work needed to solve this problem.

OpenStudy (jfraser):

same problem, different numbers

OpenStudy (anonymous):

can you do it and give me the final answer

OpenStudy (jfraser):

no, but i'll tell you if your answer is correct ;-)

OpenStudy (anonymous):

109 A°C 382.15 K

OpenStudy (jfraser):

too high

OpenStudy (anonymous):

imm going to do it again and show all work, can you check it?

OpenStudy (jfraser):

sure

OpenStudy (anonymous):

hbow do i get the mass given?

OpenStudy (anonymous):

molar mass is 58.6, correct?

OpenStudy (jfraser):

you don't need to use molar amss, because it already tells you the moles of NaCl. plug them in as is

OpenStudy (anonymous):

would it be 234

OpenStudy (jfraser):

4 moles of NaCL would weigh 234g, but you don't use grams in these equations, you use moles. it tells you 4 moles of NaCl, so plug in 4 moles of NaCl

OpenStudy (anonymous):

4 moles of solute* .51C / 2.05 Kg ?

OpenStudy (anonymous):

.995 C

OpenStudy (anonymous):

274.145 K

OpenStudy (jfraser):

that's the change in the boiling point, so the actual boiling point will be?

OpenStudy (anonymous):

how do you do that subtract the two?

OpenStudy (anonymous):

274.145- 273.66?

OpenStudy (jfraser):

what temp does water normally boil at?

OpenStudy (jfraser):

what temp does water normally boil at?

OpenStudy (anonymous):

212 373.15 k

OpenStudy (anonymous):

373.15-274.145

OpenStudy (jfraser):

the change in temp is given in C, not F or K, so keep your units straight and this will all make a whole lot more sense. water boils at 100C. the CHANGE in temp of the BP is 0.99C the NEW BP of this solution will be the sum of the two, because it's always an elevation in the boiling point

OpenStudy (anonymous):

so boiling point changes by 2 degrees

OpenStudy (jfraser):

the BP changes by 1 degree. on what planet does 0.99 = 2?

OpenStudy (anonymous):

lol thats what i meant but i hit 2

OpenStudy (jfraser):

ok then

OpenStudy (anonymous):

doesnt it go down?

OpenStudy (jfraser):

BPs are always raised by solutions, FPs are always decreased by solutions

OpenStudy (jfraser):

it has to do with the interactions that the solute molecules create in between the solvent molecules. The solute gets in the way of the phase change, no matter what direction

OpenStudy (anonymous):

.01 right